# AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications

### v2

@@ 2. Distance error of vehicles and VRUs
[126] Sensing: Camera, real time object detection
@@ 3. Distance to conflict point
Ours Sensing: Camera, YOLOv8 object detection.
@@
Computing: Server end GPU
Develops a DT pipeline to demonstrate use cases (including intersection safety warning and ATSC) in urban settings. TTC 1. Accuracy for the warning issued
@@ 2. Granular latency per frame
Intersections, where sixty percent of crashes happen [135, 136], are critical bottlenecks of an urban transportation network. To improve urban road safety and increase traffic capacities, safety warning is the key. Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application. The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency (see Fig. 3).
Resource optimization
Due to the computational needs of complex computer vision models and the high volume of video data, ensuring scalability in a video analytics pipeline requires optimization of its configuration to maintain performance. However, the system’s performance is impacted by factors such as the video content, network conditions, and available computational resources. As a result, it is crucial to implement a dynamic, real-time optimization mechanism that adjusts key configuration parameters—such as resolution, frame rate, and bitrate—based on these varying conditions. This adaptive approach allows the system to continuously balance performance with resource efficiency, ensuring scalable and reliable video analytics. Accordingly, we have equipped our video analytics pipeline with a missingResource Optimization (see Fig. 2) component that continuously adapts the system’s configuration parameters to maintain its performance.
In addition to dynamic adaptation of the system’s configuration, we can reduce latency and GPU consumption by identifying Regions of Interest (RoIs) where pedestrians might be in danger, using lightweight processing (e.g., low resolution, small models). As shown in Table VIII, smaller models (“YOLOv8s” in Column 4) incur significantly lower latency. Larger models (“YOLOv8x” in Column 6) with higher resolution and frame rate are only triggered when a critical danger area, i.e., RoI, is detected. For example, detecting large objects such as vehicles does not require large models or high resolution. Therefore, we can use a smaller model and lower resolution to detect vehicles and their trajectories, and based on that, determine the danger areas where pedestrians may be at risk or in the blind zones of vehicles. These identified danger areas are then processed using larger models and higher resolution to detect pedestrians at risk and notify them or the vehicles if necessary. In our intersection safety warning system (shown in Fig. 3), we have embedded an RoI Specifier element to facilitate this approach.
AI models
A variety of AI models are developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) [137], which encodes road user trajectories into quantized latent codes to account for heterogeneity in road users, PI-NeuGODE (Physics-Informed Graph Neural Ordinary Differential Equations), using physics-informed graph neural ordinary differential equations [138, 139], and uncertainty quantification (UQ) to characterize the predictive confidence [140]. These probabilistic trajectory prediction methods enhance pedestrian safety at intersections by incorporating UQ into risk assessment. The Kalman Filter (KF) [141] performs prediction by recursively estimating the state of a moving object using a linear dynamic model with Gaussian noise. Trajectron++ [142] models multimodal future trajectories by combining recurrent neural networks with conditional variational autoencoders, enabling probabilistic and socially-aware predictions. We present the predicted trajectories and performance comparison in Fig. 7, where the performance is evaluated using Average Displacement Error (ADE, i.e., the mean Euclidean distance between predicted and ground-truth positions over all future time steps) and Final Displacement Error (FDE, i.e., the Euclidean distance at the final prediction step).
Method ADE $\downarrow$ FDE $\downarrow$
Kalman Filter 0.91 1.92
Trajectron++ 0.77 1.34
ADE and FDE comparison. Lower values are better ($\downarrow$).
Figure 7: Performance comparison for trajectory prediction
Evaluation
In safety-critical applications, it is infeasible to perform field experiments for such assessment. DT simulation is thus used to assess the effectiveness of the system.
a. End-to-end latency
Measuring latency could be particularly challenging due to the need for cross-network timing synchronization of devices and compute server. Table VIII presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as shown in Fig. 3. We measured the latency for three sizes of YOLOv8 object detection model: YOLOv8s (small, $\sim$11.2 million parameters), YOLOv8m (medium, $\sim$25.9 million parameters), and YOLOv8x (large, $\sim$68.2 million parameters). Larger models offer improved detection accuracy but come at the cost of increased latency and higher GPU and memory usage [143]. The pipeline has been systematically optimized in terms of memory and resource usage. These elements usually run on an edge device or server. The results identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization.
TABLE VIII: Average latency and Standard Deviation for different pipeline elements.
Metric Inference Elements Downstream Elements
Reception Pre-processing Object Detection Model Object Tracking MQTT Msg Creator MQTT Msg Retrieval
Small (YOLOv8s) Medium (YOLOv8m) Large (YOLOv8x) Ethernet Wi-Fi LTE 5G
Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 4.034 (0.084) 7.216 (0.086) 11.140 (1.800) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
b. Safety message generation accuracy
Figure 8: A simulated pedestrian in CARLA (left) surrounded by cars with real-world positions (right).
To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $10$ minutes and generating a total of $232$ virtual pedestrians in CARLA. We then compared the number of collision warning messages with the actual number of simulated collisions. The process is illustrated in Fig. 8.
(a) TTC (in seconds)
(b) Distance (in pixels)
Figure 9: ROC curves for trajectory prediction to determine optimal thresholds
Note that metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point). TTC is commonly used, while PET is typically for post-event analysis. To select the optimal thresholds of TTC, we first generate the ROC (Receiver Operating Characteristic) curve in Fig. 9. Fig. 9(a) illustrates the relationship between the true positive rate (TPR) and the false positive rate (FPR) across various TTC threshold values, ranging from 0.1 to 1.2 seconds. The optimal TTC threshold is 1.1 seconds (indicated by a blue square), at which the TPR reaches $0.958$ and the FPR is $0.4$. To compute TTC, we need to compare each pair of predicted trajectory points in the next $t$ time steps, determined by a danger distance threshold that is defined as when the proximity between a vehicle and a pedestrian constitutes a dangerous interaction. Fig. 9(b) presents the TPR and FPR for the threshold of danger distances ranging from $5$ to $100$ pixels. The optimal threshold is valued at $30$ pixels, which yields a TPR of $0.958$ and an FPR of $0.345$. Based on the above two thresholds, we run the collision prediction model. The resultant confusion matrix for predicted collisions under the selected thresholds is given by $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]]$.
c. Human response time assessment
Figure 10: Response time determination in the VR experiment
Figure 11: Survey results from the VR experiment
Would issuing warnings to pedestrians help reduce users’ response time and increase their safety awareness? To test this hypothesis, we designed virtual reality (VR) experiments in Unity3D [144], where warnings are provided via voice and text displayed on the VR headset Meta Quest 3. There are two traffic scenarios, one involving an interaction between the participant (i.e., a pedestrian) and an oncoming scooter, and the other between the participant and an oncoming vehicle. In each scenario, participants first receive no warning and then a warning, and their response times are recorded. As illustrated in Fig. 10, when traffic approaches at time $t$, the user either receives a warning or does not. The user stops after a response time $t_{r}$, which is then recorded. The response time distributions are presented in Fig. 12, with the ‘no warning’ condition shown in yellow and the ‘warning’ condition in blue. The issuance of warning has reduced people’s average response time by $0.62\mathrm{s}$ (in Scenario 1) and $1.11\mathrm{s}$ (in Scenario 2), respectively. The demographics of the participants are shown in the upper section of Fig. 11. The bar chart summarizes the participants’ survey responses regarding VR and traffic realism, the perceived importance of safety warnings, and comfort within the VR environment, with mean and standard deviation indicated on the top left corners of each subfigure.
Figure 12: Comparison of response times with and without a warning. A scooter (in Scenario 1) and a vehicle (in Scenario 2) poses the danger, respectively.
IV-C3 Discussions
We would like to highlight that our NYC’s testbed deployment demonstrates generalizability. As the biggest metropolis of the U.S., NYC’s dense population, limited space, and multimodal transportation systems (consisting of public transit with buses and subway, driving, ferry) motivate us to study the interactions between automobiles and VRUs (including pedestrians, cyclists, and surging micromoblity and e-mobility). Located in one of the world’s busiest regions - Manhattan of the NYC, the COSMOS testbed offers a natural laboratory for us to collect abundant data and conduct live experiments, which would otherwise not be possible.
We leverage the COSMOS testbed to design and implement sensor (e.g., camera, LiDAR) data analytics systems and evaluate them in urban settings to ensure they can scale to thousands of connected intersections. To achieve this, we ensure that the designed analytics systems are efficient in resource consumption, particularly network bandwidth and GPU usage. We also ensure that these systems are modular and deployable in a distributed manner across multiple edge servers and cloud infrastructure. Fig. 3 demonstrates such a modular and distributable system. This design enables flexibility, allowing the system to be deployed in settings where a single edge server cannot handle the full workload, requiring distribution across several edge servers. Therefore, the system must be designed such that each module can be easily deployed on a different server and can seamlessly communicate with other modules, regardless of whether they reside on the same server. Additionally, the AI models used must be replaceable or generalizable to accommodate different environments, camera angles, heights, and other variations. With these principles in mind, we use the COSMOS testbed to develop scalable systems.
The NYC’s dense high-rise buildings enable us to leverage legacy infrastructure for traffic sensing and monitoring, and facilitates the instrumentation of cameras and communication nodes. Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 146], which requires customized infrastructure and user-end devices. For rapid and scalable deployment, our pilot experiments utilize contemporary protocols, which hold the potential to further leverage the city’s legacy traffic cameras and wireless communication infrastructure for scalability and cost-effectiveness. Our technologies could be replaced by emerging methods that can achieve lower latency and higher bandwidths. For instance, software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications. Since major intersections in the neighborhood has similar sensor suites, it would not need additional infrastructure deployment to generalize our system to network wide multi-intersection settings.
We will extend our use cases to multi-pedestrian warning scenarios. Smartphone‑to‑cloud beaconing with an Adaptive Multi‑Mode (AMM) scheme enables real‑time location sharing from multiple devices, allowing the server to issue timely risk alerts [147]. A Vehicle‑to‑Pedestrian (V2P) system [148] leveraging Bluetooth Low Energy (BLE) advertising on smartphones broadcasts standardized Personal Safety Messages (PSM) to address the multi‑device communication challenge in road safety. Moreover, efficient multi-device communication is essential for real-time data sharing and coordination among mobile users in smart urban environments. Key challenges include ensuring reliable connectivity in dense settings, minimizing communication latency, and maintaining user privacy. Advanced techniques such as device-to-device (D2D) communication [149], opportunistic networking [150], and privacy-preserving data aggregation [151] have been developed to address these issues.
V Conclusions and open questions
In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management. Then, leveraging the unique characteristics of NYC, we propose a DT architecture and present a safety-critical use case, namely, intersection safety warning. Three evaluation methods are performed, including measuring latency, simulating collision reduction, and VR experiments for human responses.
@@ V Conclusions and open questions
A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for uban traffic management. Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
V-A Emerging trends
@@ III. Integrated sensing and communication
V-A2 Emerging DT applications
With the increasing demand for urban passenger and goods delivery, emerging technology like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space. DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
@@ V-A2 Emerging DT applications
Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC). With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes. There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [169, 170, 171, 172, 173, 174, 175, 176, 177]. With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [178, 179, 180, 164, 181, 182].
Driven by growing urban populations, low-altitude economy aims to exploit vertical space in urban environments, advanced by drones and eVTOL (electric vertical take-off and landing). It will foster wide applications, ranging from urban air mobility, urban logistics, agriculture, emergency service, to infrastructure inspection, surveillance and security. Critical challenges arise in sensing, modeling, predicting urban aerodynamics, and collision avoidance with high-rise buildings and flying objects. Open questions include market demand estimation, infrastructure planning, operation and design. DTs hold the potential to the adoption of these emerging technologies, for its power in simulating emerging demand patterns, modeling their kinetics and dynamics, routing and charging behaviors. To build such DTs, CPS technologies and AI methodologies would facilitate precise sensing and perception [146], reliable and fast communication and networking [183], seamless multimodal coordination (such as drone-truck-mobile robot delivery) [184], as well as safe and optimal decisions [185].
V-A3 AI models
@@
The presence of human factors along the DT pipeline could complicates the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [208]. Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs. For DTs that involve the feedback to humans, human-in-the-loop test is widely used [209, 210, 208]. Recent years have witnessed a growing trend of using augmented reality and virtual reality to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [134, 211]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [212], could help test the performance of some physical components “live,” while allowing the rest to be simulated within a DT. Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
Benchmark datasets and methods
To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use. Instead, we hope that an AI-powered DT could be generalized to diverse tasks, transferred to diverse spatiotemporal settings, and shared for co-development among global researchers. Accordingly, we need to standardize application scenarios, benchmark datasets and methods, and unified test environments, for repetitive training and test [213]. Benchmark datasets and methods necessitate performance comparison of any newly proposed AI methods against the state-of-the-art (SOTA) methods. Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability. For example, standard test platforms, such as Gym [214], Flower [215], PettingZoo [216, 217], and Ray RLlib [218], are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
@@
Moreover, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue to address this gap [219, 220, 221].
V-B Challenges and open questions
V-B1 Closing real-to-sim-to-real gap

### v3

@@ 2. Distance error of vehicles and VRUs
[65] Sensing: Camera, real-time object detection
@@ 3. Distance to conflict point
[69] Sensing: Camera, YOLOv8 object detection.
@@
Computing: Server end GPU, resource allocation
Develops a DT pipeline to demonstrate use cases (including intersection safety warning) in urban settings. Distance to the danger zone 1. Localization accuracy
@@ 2. Granular latency per frame
4.2 Adaptive traffic signal control
Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC). With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes. There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [95, 39, 104, 97, 264, 146, 112, 34, 79]. With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [243, 86, 15, 62, 56, 127]. Related work are summarized in Tab. 7. The last column indicates whether the optimization accounts for the safety of VRUs or not.
Table 7: Summary of papers using (F)RL for ATSC
Ref. RL FL DT Simulator Demand Network VRU
[95] DDPG SUMO
[39] SUMO
[104] A2C Vissim
[97] Vissim
[264] SUMO
[62] A3C FedAvg SUMO
[243] DQN FedAvg Cityflow
[86] PPO FedAvg SUMO
[15] DQN FedAvg SUMO
[146] A2C SUMO
[112] A2C SUMO
[34] A2C SUMO
[79] PPO SUMO
[167] PPO SUMO
[173] D3QN SUMO
[241] DDQN SUMO
Demand: Real-world traffic demand data; Network: Real-world road network.
FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
5 Physical Testbeds for DT validation
Physical testbeds are essential for verification, validation, and testing of DTs in diverse and out-of-distribution scenarios. It offers a platform for sensor instrumentation, data collection, actuation, to close the information flow and control feedback loop. Datasets collected from a testbed, if open-sourced, could be used by other researchers who cannot afford building a testbed. Below, we summarize existing testbeds in Tab. 8 based on their types, namely, real-world, closed-track, or scaled-down. Except COSMOS focusing on the dense urban traffic environment, these testbeds are primarily developed for the testing of vehicular technologies and control.
Table 8: Summary of real-world and scaled down testbeds in the U.S.
Type Testbed Open Data Software Hardware Testing scenarios
Real-world COSMOS (NY) [171] software-defined networking (SDN), edge cloud, radio software stack radio access & wireless nodes, mmWave antennas, cameras, GPU/FPGA accelerators V2X networking; traffic sensing; edge computing experiments
Toyota Testbed (CA) [215] AWS cloud platform, traffic simulation platform edge computing, vehicle on-board sensors, full-scale vehicle, onboard computing units, CPU/GPU nodes CAV platooning, ATSC algorithm evaluation, mixed traffic, vehicle control testing
SMTG (NJ) [92] computing cluster, Carla/SUMO co-simulation, Unity visualization lidar, camera, Bluetooth, NVIDIA Jetson AGX Xavier on edge, RSU, Cellular modems real world traffic detection, V2X, DT evaluation, VRU safety.
STARlab (WA) [210] STAR Detection App, VISSIM simulation platform, edge station software V2X Edge Node, Bluetooth and WiFi radios, Raspberry Pi V2X safety, mobility detection
GoMentum (CA) [35] HD Map Database, V2X middleware, sensor middleware full-scale AVs, edge computing, lidar camera, RTK-GPS signalized intersection testing, Construction zone testing, VRU testing
Purdue Testbed (IN) [263] autonomy stack, ROS, cloud-based VLM, whisper, Autoware full-scale AV, lidar, radar, camera, GNSS, high bandwidth networking highway driving, intersection safety, VLM-controlled AV, parking lot testing
Closed track Mcity (MI) [163] unreal virtual Mcity, V2X software stack, driving simulation 32-acre closed test track, perception sensors, RSUs, DSRC, Roadside LiDAR, full-scale AVs ATSC, CAV testing, signalized intersections, AV perception/decision experiments
Scaled down Duckietown (MA) [162] Robot Operating System (ROS) server, signal control framework scaled autonomous vehicle and road, traffic light, Wi-Fi lane following, city navigation, traffic light control
F1Tenth (MA,PA,VA) [14] ROS, Gazebo simulation, RViz visualization, oracle node 1/10 autonomous racecar, NVIDIA Jetson TX2, Wi-Fi, UDP sim-to-real, vehicle racing, vehicle control
Ccity (NY) [168] ROS, self-built simulation, driving emulator 1/8 vehicle with Jetson Orin Nano, camera, UWB localization, traffic signal, Wi-Fi, Raspberry Pi AV lane following, platooning, AV control, traffic signal control, human driving behaviour testing
IDS3C (NY) [26] UDP sockets, IDS 3D city, ROS, Unity engine, Microsoft AirSim plugin 1/25 robotic ground cars, Raspberry Pi, Wi-Fi, UDP, VICON motion capture system CAV control, CPS validation, eco driving, shared mobility, sim-to-real
6 Conclusions and open questions
@@
In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management. Then, how DTs are leveraged for use cases including intersection safety warning and traffic signal control are introduced. Following that, the testbeds for DT testing and validation are presented. A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for urban traffic management. Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
6.1 Emerging trends
@@ III. Integrated sensing and communication
6.1.2 Emerging DT applications
With the increasing demand for urban passenger and goods delivery, emerging technologies like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space. DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
@@ 6.1.2 Emerging DT applications
6.1.3 AI models
@@
The presence of human factors along the DT pipeline could complicate the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [42]. Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs. For DTs that involve the feedback to humans, human-in-the-loop test is widely used [244, 113, 42]. Recent years have witnessed a growing trend of using augmented reality and virtual reality to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [214, 105]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [231], could help test the performance of some physical components “live,” while allowing the rest to be simulated within a DT. Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
Benchmark datasets & shared testbed resources
@@
Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems. However, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue to address this gap [129, 144, 30].
To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use. Instead, we hope that an AI-powered DT could be generalized to diverse tasks, transferred to diverse spatiotemporal settings, and shared for co-development among global researchers. Accordingly, we need to standardize application scenarios, benchmark datasets and methods, and unified test environments, for repetitive training and test [43]. Benchmark datasets and methods necessitate performance comparison of any newly proposed AI methods against the state-of-the-art (SOTA) methods. Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability. For example, standard test platforms, such as Gym, Flower, PettingZoo, and Ray RLlib, are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
While data and methods can often be shared, advancing DTs ultimately depends on shared physical testbeds. These real-world platforms are essential for validation and learning, yet remain costly and space-constrained—calling for new, collaborative approaches from the community. Recent years have seen a growing effort to crowd-source existing testbeds pushed by both agencies and researchers. For example, augmented from the Mcity Test Facility by remote access, Mcity 2.0 [150] enables researchers to connect to the site via a secure cloud interface, run experiments without being physically present, and interact with a mixed-reality environment that overlays virtual actors onto the physical track. In this way, Mcity 2.0 operationalizes standardized, repeatable testing by tightly integrating simulation, physical infrastructure, and remote vehicle control in one unified platform.
6.2 Challenges and open questions
6.2.1 Closing real-to-sim-to-real gap